# Artificial-Intelligence Generated Code Considered Harmful: A Road Map for Secure and High-Quality Code Generation

### v1

@@ 1 Introduction
The security implications of AI code generation are largely ignored by the industry: Copilot Voice has been introduced as a new feature that allows inexperienced developers to generate full programs by simply speaking to the AI assistant, marketed as a new way to “write code without the keyboard” [16]. Naïve trust in AI code generation can lead to significant security vulnerabilities and degradation in code quality. Understanding the traits and limitations of the LLM-generated code is important to facilitate the adoption of LLMs in future software engineering practice. In October 2023, the White House issues an Executive Order on “Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence”, which emphasize the importance of security in such AI use cases [10, 4].
@@
We also found that the security and quality of LLM-generated code is lacking, compared to its human-written equivalent. The security issues found by Clang static analyzer in LLM-generated code are consistently higher than human-written code: by 11.2% in LeetCode and 7.1% in algorithm tasks. Additionally, LLM-generated code is 1.19x more complex than human-written code in LeetCode and 1.26x in algorithm tasks.
A feedback loop is a state-of-the-art practice carried out by various research groups [32, 47, 36] to improve LLM’s code generation. Therefore, our study also investigates the effectiveness of a feedback loop in improving the security and correctness of the generated code by iteratively feeding the results of the evaluation back to the LLM model. We show that when we asked the model to fix the discovered security issues, it can introduce new bugs in code that was previously bug-free; moreover, the re-generated version has higher complexity per line. We also observe that by requesting shorter line lengths, GPT-4o yields subsequent code with 11.3% lower cyclomatic complexity per line, and surprisingly, it also fixed 17.9% security issues found by Clang static analyzer. In contrast, by requesting fewer lines of code, GPT-4o yields code with 18.4% higher complexity per line, and fixed 43.2% of security issues.
During the course of our study, we have identified 116 security issues in public GitHub repositories, which we have responsibly disclosed to the repository owners. For the software security issues of the code generated by OpenAI’s GPT-4o model, we have reached out to OpenAI for their further improvement. We will open source EXACT and our test data. We discuss the full details in the Ethics and Open Science Compliance appendix.
@@ 2.2 Automated Code Improvement
LLMs have been using Reinforcement learning (RL) for automated code improvement. RL updates a model’s parameters through interactions with human feedback or the environment, and has shown to be effective in improving various AI models’ performance [32, 47]. Reinforcement learning from human feedback (RLHF) uses human feedback to fine tune a model’s parameters [30]. Despite the fact that RLHF can be used to improve the performance of LLM code generation, scalable deployment in practice has been limited by the capability of human evaluators to provide meaningful feedback to the outputs [36]. To cope with the lack of human feedback, OpenAI has proposed a fine-tuned model based on GPT 4, namely CriticGPT, to automatically critique the LLM-generated code, where the feedback is used to improve the subsequent generations  [36]. A pool of bugs and human feedback originated from OpenAI’s previous RLHF pipeline is used as the training set of the fine-tuning of CriticGPT [36]. Likewise, GitHub Copilot has proposed a secondary LLM model to “approximate the behavior of static analysis tools” to provide feedback to the outputs of the primary LLM model  [6]. Since these systems have been integrated into the LLM pipeline, our work aims to investigate the security of the overall system that may include these blackbox components.
@@ 3.1 Sample Task 1: Finding a Buffer’s Size
As shown in the Table 1, GPT-4o has a high success rate (but not 100%) in filling in the buffer size when the buffer size is a constant integer. When the buffer size requires calculation, the success rate drops significantly. The success rate is the lowest when the buffer size is calculated using floating-point numbers, possibly due to the model’s lack of understanding of float to integer casting. To our surprise, GPT-4o’s success rate drops from 92.8% to 41.6% when the buffer size is calculated from the length of alphabet with a missing letter, showing that the model concludes the buffer size based on the pattern of the alphabet, rather than the actual length of the string. Figure 2 shows the distribution of the buffer size filled in by GPT-4o in the missing letter example over 1000 trials. We were also surprised that when we give the model a constant integer buffer size, but at the same time, present a irrelevant calculation of the length of a string, the success rate drops from 99.3% to 31.2%. In rare cases (0.08% of all results), the model has even filled the size (an expected int value) with meaningless non-English characters and symbols. This preliminary study demonstrates that GPT-4o has a high chance of generating memory bugs in C code, and motivates us to further investigate the security of LLM-generated code.
@@ 3.3 Sample Task 3: Implementing Graph Data Structure
This task entailed generating a simple graph implementation, with functions for creating (allocating) a graph, adding and removing edges, printing the graph, and freeing (deallocating) the graph memory from the heap. The relevant allocation code – creating a graph of size V – is shown in Figure 3. Note that the human-developed code, shown on the left, has checks for the value of V, whereas the LLM-generated code, on the right, does not contain such checks. The LLM code has two issues. First, it does not check the result of malloc() for NULL, hence opening the first entry for potential NULL pointer dereferences if the memory allocation fails. Second, the LLM-generated code does not check for a potential negative value of V. If a negative V is passed as an argument, the call to malloc() on line 5 will most likely result in a NULL pointer (since malloc() takes an unsigned argument, the argument is interpreted as a positive value at the size of size_t_max - abs(V) + 1, where size_t_max is the max value of the type), or, less likely, in a very large memory allocation. Therefore, memory allocation either fails or its size is in the control of a potential adversary. We refer to this as malloc overflow, as a term used by Clang analyzer.
@@ 3.4 Motivations
Aligned with the national goal to manage the risk of generative AI [10, 4], our research investigate the security of LLM-generated code. Our research is motivated by the following questions:
@@ 4.1 LLM Code Generation
We use OpenAI’s GPT-4o model to generate the full task solution for each coding task. As shown in Figure 4, for each coding task, we provide its task description as prompt to GPT-4o API, and collect the coding outputs.
@@ 4.2.2 Data Structures and Algorithms
Unit testing. Test cases are manually written for unit testing with the help of CUnit [9]. We examine the human-written code to understand the functionality provided by the data structure or algorithm, such as insertion, deletion, sorting, etc. We then write test cases, divided into 2 categories: (1) regular cases, and (2) edge cases. Regular cases are typical scenarios that a function is expected to handle, such as inserting positive integers into a linked list, perform a merge sort on an unsorted array, etc. Edge cases are crafted in a way to test a function’s protection against unexpected inputs, such as dequeuing from an empty queue, performing a breadth-first search on an empty graph, sorting an empty array, etc. We ran the same test cases on both GPT-4o-generated and human-written code, since we prompt GPT-4o to implement the data structures and algorithms with the same header files or function declarations as used in the human-written code. By testing GPT-4o-generated and human-written code with the same test cases, we verify that they are achieving the same functionalities.
Fuzzing. In addition to unit testing, we use AFL [2] to fuzz both GPT-4o-generated code and human-written code. We create an entry point to fuzz each data structure and algorithm. This entry point is the main method in each program where it reads an input file that contains various instructions and input values. For example, the line insert 10 in an input file will carry out the insert operation in the binary search tree and insert the value 10 into the tree. The input file is used as a seed in fuzzing. Each data structure and algorithm has its own set of instructions to make sure that all implemented functionalities will be tested. Since we only focus on functionality validation, any invalid instructions mutated by AFL [2] in the process of fuzzing will be ignored in the entry point. However, values will be mutated during fuzzing and GPT-4o-generated code and human-written code need to handle any kinds of invalid inputs. As it was not relevant to the comparison, the entry point code was not included in the security and complexity analysis.
@@ 4.3.1 Security
We use the Clang static analyzer [1] to perform static analysis on both GPT-4o-generated and human-written code. Clang provides a set of default checkers, such as null dereference, memory leak, and null arguments. In addition, Clang provides a range of experimental (advanced) checkers in several categories, as follows. “Core” experimental checkers include detectors for pointer arithmetic, invalid casting/conversion, etc. “Security” experimental checkers look for errors such as array index out of bounds, malloc overflow, etc. “Unix” experimental checkers look for issues such as memory leaks or null pointers passed to string functions. While the experimental checkers might, in theory, emit a higher number of false positives, our manual validation of the reported errors indicate a negligibly low rate of false positives.
@@ 5.2 Security Analysis
We perform static analysis on both GPT-4o-generated and human-written code using Clang static analyzer [1]. Table 4 shows the analysis results of 220 files (202 LeetCode problems, 18 data structures and algorithms). Overall, GPT-4o-generated code has 10.3% more security issues than human-written code. The three categories of issues that have relatively higher count are malloc overflow, array index out of bounds, and memory leak. We found that GPT-4o-generated code has 32.8% more malloc overflow issues than human-written code. This is a consequence of GPT-4o’s tendency of assuming all inputs are valid (Figure 3). The counts for array index out of bounds issues are similar for both GPT-4o and human. Figure 5 illustrates the similar mistakes made by both GPT-4o and human in one of the LeetCode problems where both code may access elements that are out of bounds. Memory leak issues, where human code has a 50% higher count than GPT-4o, are mainly due to the problem description, where the LeetCode problems specifically mention that the caller will invoke free(). However, the memory leak issues do exist if the code snippet are used outside the context of these LeetCode questions.
Other than issues reported by Clang, we also look at the implementation details between GPT-4o-generated and human written code on data structures and algorithms. We do not focus on the implementation details of LeetCode solutions because the main objective of a LeetCode problem is to provide a solution that solves a very particular problem with a given set of constraints that passes all the test cases provided.
We found that GPT-4o has a tendency of generating exact same code as human-written code. We will refer to this as “code parroting” for the rest of the paper. Note that we do not include any implementation details of human-written code in the prompts with the exception of header files or function declarations. In GPT-4o-generated code, 2 out of the 18 files in data structures and algorithms are the exact replica of human written code, as follows. In AVL tree’s implementation, GPT-4o reproduces the human-written code from GitHub, including the helper functions. Similarly, in the graph data structure implementation, GPT-4o reproduces the human-written code, but fails to add validation for input values (Figure 3). Code parroting is extremely harmful considering the fact that malicious code could possibly be the training data of LLM (more commonly known as data poisoning if malicious information is injected into training data on purpose). Harmful code will then be generated by LLM and possibly utilized by developers.
@@ 5.3 Complexity Analysis
SCC [25] is the tool that we utilize to measure cyclomatic complexity and lines of code (LoC, excluding comments and blank lines) per file. We calculate mean, median, geometric mean, and trimmed mean for both complexity and normalized complexity (complexity per LoC).
As shown in Table 5, GPT-4o-generated code seems to have lower complexity on the surface. However, it would be incorrect to assume that this lower complexity indicates high-quality code. First, low (absolute) complexity is an artifact of GPT-4o generating less code than the human equivalent for the same task. In fact, human-written code always has a lower normalized complexity, which is a more indicative metric of the efforts needed to understand and maintain a typical snippet of code in a given file. Second, the low complexity is due to GPT-4o generating “bare-bones” code. For instance, the GPT-4o-generated hash map implementation (Figure 6) uses a simple hash function that does not check for collisions, while the human-written implementation has a more sophisticated hash function that checks for collisions (Figure 7). This example illustrates one of the reasons why the GPT-4o-generated code has lower cyclomatic complexity (when computing complexity for the entire file, i.e., without normalizing for lines of code). Another example is in Figure 8 where the human-written implementation validates input values before carrying out certain operations in one of the LeetCode problems, whereas GPT-4o makes the assumptions of all inputs are valid. This undoubtedly increases the cyclomatic complexity of human-written code but illustrates that human-written code uses defensive programming, which in turn mitigates various security threats.
@@ 6.1 Rationale and Setup
State-of-the art practice suggests using a reinforcement loop (Section 2.2) to improve code quality when using LLMs. Therefore, we constructed a feedback loop as a way to improve GPT-4o-generated code, starting by trying to eliminate the security issues identified in Section 5.2. For this experiment, we included 60 LeetCode files (solutions) generated by GPT-4o, as follows: 30 files with security issues (found by Clang), and 30 files without security issues. The list of security issues is listed in Table 7. We ask GPT-4o to regenerate these files based on different security categories. In each loop, prompts for files with security issues will be accompanied by prompts for files that have no security issues. For instance, we specifically ask GPT-4o to fix only malloc overflow issues in this loop with a total a 18 prompts related to files that have malloc overflow issues along with another 18 prompts related to files that have no malloc overflow issues. Note that the prompts (LeetCode problem descriptions) used in the loops are the same as the non-feedback setup, but with additional instructions that ask GPT-4o to make sure there are no certain security issues in the code. Since there are 3 categories of security issues, a total of 3 loops will be conducted, each loop to fix a specific security issue.
@@ 6.2 Results
Files with security issues. For files that did contain security issues, GPT-4o manages to lower the number of security issues found after one iteration on malloc overflow and null dereference from 40 to 24 and 3 to 2, respectively, as shown in Table 8. However, the security issues are not completely removed, even though we specifically asked GPT-4o to ensure the code will not have these security issues. Furthermore, for array index out of bounds issues, GPT-4o generated 3 more such issues. This is a critical finding since it shows the possibility that LLM may generate code that is the exact opposite of what users have specified in the prompt.

### v2

@@ 1 Introduction
The security implications of AI code generation are largely ignored by the industry: Copilot Voice has been introduced as a new feature that allows inexperienced developers to generate full programs by simply speaking to the AI assistant, marketed as a new way to “write code without the keyboard” [16]. Naïve trust in AI code generation can lead to significant security vulnerabilities and degradation in code quality. Understanding the traits and limitations of the LLM-generated code is important to facilitate the adoption of LLMs in future software engineering practice. In October 2023, the White House issued an Executive Order on “Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence”, which emphasize the importance of security in such AI use cases [10, 4].
@@
We also found that the security and quality of LLM-generated code are lacking, compared to its human-written equivalent. The security issues found by Clang static analyzer in LLM-generated code are consistently higher than human-written code: by 11.2% in LeetCode and 7.1% in algorithm tasks. Additionally, LLM-generated code is 1.19x more complex than human-written code in LeetCode and 1.26x in algorithm tasks.
A feedback loop is a state-of-the-art practice carried out by various research groups [32, 47, 36] to improve LLM’s code generation. Therefore, our study also investigates the effectiveness of a feedback loop in improving the security and correctness of the generated code by iteratively feeding the results of the evaluation back to the LLM model. We show that when we asked the model to fix the discovered security issues, it can introduce new bugs in code that was previously bug-free; moreover, the re-generated version has higher complexity per line. We also observe that by requesting shorter line lengths, GPT-4o yields subsequent code with 11.3% lower cyclomatic complexity per line, and surprisingly, it also fixed 17.9% of the security issues found by Clang static analyzer. In contrast, by requesting fewer lines of code, GPT-4o yields code with 18.4% higher complexity per line, and fixed 43.2% of security issues.
During the course of our study, we have identified 116 security issues in public GitHub repositories, which we have responsibly disclosed to the repository owners. For the software security issues of the code generated by OpenAI’s GPT-4o model, we have reached out to OpenAI for their further improvement. We will open source EXACT and our test data.
@@ 2.2 Automated Code Improvement
LLMs have been using Reinforcement learning (RL) for automated code improvement. RL updates a model’s parameters through interactions with human feedback or the environment, and has been shown to be effective in improving various AI models’ performance [32, 47]. Reinforcement learning from human feedback (RLHF) uses human feedback to fine-tune a model’s parameters [30]. Despite the fact that RLHF can be used to improve the performance of LLM code generation, scalable deployment in practice has been limited by the capability of human evaluators to provide meaningful feedback to the outputs [36]. To cope with the lack of human feedback, OpenAI has proposed a fine-tuned model based on GPT 4, namely CriticGPT, to automatically critique the LLM-generated code, where the feedback is used to improve the subsequent generations  [36]. A pool of bugs and human feedback originated from OpenAI’s previous RLHF pipeline is used as the training set for the fine-tuning of CriticGPT [36]. Likewise, GitHub Copilot has proposed a secondary LLM model to “approximate the behavior of static analysis tools” to provide feedback to the outputs of the primary LLM model  [6]. Since these systems have been integrated into the LLM pipeline, our work aims to investigate the security of the overall system that may include these blackbox components.
@@ 3.1 Sample Task 1: Finding a Buffer’s Size
As shown in Table 1, GPT-4o has a high success rate (but not 100%) in filling in the buffer size when the buffer size is a constant integer. When the buffer size requires calculation, the success rate drops significantly. The success rate is the lowest when the buffer size is calculated using floating-point numbers, possibly due to the model’s lack of understanding of float to integer casting. To our surprise, GPT-4o’s success rate drops from 92.8% to 41.6% when the buffer size is calculated from the length of alphabet with a missing letter, showing that the model concludes the buffer size based on the pattern of the alphabet, rather than the actual length of the string. Figure 2 shows the distribution of the buffer size filled in by GPT-4o in the missing letter example over 1000 trials. We were also surprised that when we give the model a constant integer buffer size, but at the same time, present an irrelevant calculation of the length of a string, the success rate drops from 99.3% to 31.2%. In rare cases (0.08% of all results), the model has even filled the size (an expected int value) with meaningless non-English characters and symbols. This preliminary study demonstrates that GPT-4o has a high chance of generating memory bugs in C code, and motivates us to further investigate the security of LLM-generated code.
@@ 3.3 Sample Task 3: Implementing Graph Data Structure
This task entailed generating a simple graph implementation, with functions for creating (allocating) a graph, adding and removing edges, printing the graph, and freeing (deallocating) the graph memory from the heap. The relevant allocation code – creating a graph of size V – is shown in Figure 3. Note that the human-developed code, shown on the left, has checks for the value of V, whereas the LLM-generated code, on the right, does not contain such checks. The LLM code has two issues. First, it does not check the result of malloc() for NULL, hence opening the first entry for potential NULL pointer dereferences if the memory allocation fails. Second, the LLM-generated code does not check for a potential negative value of V. If a negative V is passed as an argument, the call to malloc() on line 5 will most likely result in a NULL pointer (since malloc() takes an unsigned argument, the argument is interpreted as a positive value at the size of size_t_max - abs(V) + 1, where size_t_max is the max value of the type), or, less likely, in a very large memory allocation. Therefore, memory allocation either fails or its size is in the control of a potential adversary. We refer to this as malloc overflow, a term used by Clang analyzer.
@@ 3.4 Motivations
Aligned with the national goal of managing the risk of generative AI [10, 4], our research investigates the security of LLM-generated code. Our research is motivated by the following questions:
@@ 4.1 LLM Code Generation
We use OpenAI’s GPT-4o model to generate the full task solution for each coding task. As shown in Figure 4, for each coding task, we provide its task description as the prompt to GPT-4o API, and collect the coding outputs.
@@ 4.2.2 Data Structures and Algorithms
Unit testing. Test cases are manually written for unit testing with the help of CUnit [9]. We examine the human-written code to understand the functionality provided by the data structure or algorithm, such as insertion, deletion, sorting, etc. We then write test cases, divided into 2 categories: (1) regular cases, and (2) edge cases. Regular cases are typical scenarios that a function is expected to handle, such as inserting positive integers into a linked list, performing a merge sort on an unsorted array, etc. Edge cases are crafted in a way to test a function’s protection against unexpected inputs, such as dequeuing from an empty queue, performing a breadth-first search on an empty graph, sorting an empty array, etc. We ran the same test cases on both GPT-4o-generated and human-written code, since we prompted GPT-4o to implement the data structures and algorithms with the same header files or function declarations as used in the human-written code. By testing GPT-4o-generated and human-written code with the same test cases, we verify that they are achieving the same functionalities.
Fuzzing. In addition to unit testing, we use AFL [2] to fuzz both GPT-4o-generated code and human-written code. We create an entry point to fuzz each data structure and algorithm. This entry point is the main method in each program where it reads an input file that contains various instructions and input values. For example, the line insert 10 in an input file will carry out the insert operation in the binary search tree and insert the value 10 into the tree. The input file is used as a seed in fuzzing. Each data structure and algorithm has its own set of instructions to make sure that all implemented functionalities will be tested. Since we only focus on functionality validation, any invalid instructions mutated by AFL [2] in the process of fuzzing will be ignored at the entry point. However, values will be mutated during fuzzing and GPT-4o-generated code and human-written code need to handle any kinds of invalid inputs. As it was not relevant to the comparison, the entry point code was not included in the security and complexity analysis.
@@ 4.3.1 Security
We use the Clang static analyzer [1] to perform static analysis on both GPT-4o-generated and human-written code. Clang provides a set of default checkers, such as null dereference, memory leak, and null arguments. In addition, Clang provides a range of experimental (advanced) checkers in several categories, as follows. “Core” experimental checkers include detectors for pointer arithmetic, invalid casting/conversion, etc. “Security” experimental checkers look for errors such as array index out of bounds, malloc overflow, etc. “Unix” experimental checkers look for issues such as memory leaks or null pointers passed to string functions. While the experimental checkers might, in theory, emit a higher number of false positives, our manual validation of the reported errors indicates a negligibly low rate of false positives.
@@ 5.2 Security Analysis
We perform static analysis on both GPT-4o-generated and human-written code using Clang static analyzer [1]. Table 4 shows the analysis results of 220 files (202 LeetCode problems, 18 data structures and algorithms). Overall, GPT-4o-generated code has 10.3% more security issues than human-written code. The three categories of issues that have relatively higher counts are malloc overflow, array index out of bounds, and memory leak. We found that GPT-4o-generated code has 32.8% more malloc overflow issues than human-written code. This is a consequence of GPT-4o’s tendency of assuming all inputs are valid (Figure 3). The counts for array index out of bounds issues are similar for both GPT-4o and human. Figure 5 illustrates the similar mistakes made by both GPT-4o and human in one of the LeetCode problems where both code may access elements that are out of bounds. Memory leak issues, where human code has a 50% higher count than GPT-4o, are mainly due to the problem description, where the LeetCode problems specifically mention that the caller will invoke free(). However, the memory leak issues do exist if the code snippet is used outside the context of these LeetCode questions.
Other than issues reported by Clang, we also look at the implementation details between GPT-4o-generated and human-written code on data structures and algorithms. We do not focus on the implementation details of LeetCode solutions because the main objective of a LeetCode problem is to provide a solution that solves a very particular problem with a given set of constraints that passes all the test cases provided.
We found that GPT-4o has a tendency of generating exact same code as human-written code. We will refer to this as “code parroting” for the rest of the paper. Note that we do not include any implementation details of human-written code in the prompts with the exception of header files or function declarations. In GPT-4o-generated code, 2 out of the 18 files in data structures and algorithms are the exact replica of human-written code, as follows. In AVL tree’s implementation, GPT-4o reproduces the human-written code from GitHub, including the helper functions. Similarly, in the graph data structure implementation, GPT-4o reproduces the human-written code, but fails to add validation for input values (Figure 3). Code parroting is extremely harmful considering the fact that malicious code could possibly be the training data of LLM (more commonly known as data poisoning if malicious information is injected into training data on purpose). Harmful code will then be generated by LLM and possibly utilized by developers.
@@ 5.3 Complexity Analysis
SCC [25] is the tool that we utilize to measure cyclomatic complexity and lines of code (LoC, excluding comments and blank lines) per file. We calculate the mean, median, geometric mean, and trimmed mean for both complexity and normalized complexity (complexity per LoC).
As shown in Table 5, GPT-4o-generated code seems to have lower complexity on the surface. However, it would be incorrect to assume that this lower complexity indicates high-quality code. First, low (absolute) complexity is an artifact of GPT-4o generating less code than the human equivalent for the same task. In fact, human-written code always has a lower normalized complexity, which is a more indicative metric of the efforts needed to understand and maintain a typical snippet of code in a given file. Second, the low complexity is due to GPT-4o generating “bare-bones” code. For instance, the GPT-4o-generated hash map implementation (Figure 6) uses a simple hash function that does not check for collisions, while the human-written implementation has a more sophisticated hash function that checks for collisions (Figure 7). This example illustrates one of the reasons why the GPT-4o-generated code has lower cyclomatic complexity (when computing complexity for the entire file, i.e., without normalizing for lines of code). Another example is in Figure 8 where the human-written implementation validates input values before carrying out certain operations in one of the LeetCode problems, whereas GPT-4o makes the assumptions that all inputs are valid. This undoubtedly increases the cyclomatic complexity of human-written code but illustrates that human-written code uses defensive programming, which in turn mitigates various security threats.
@@ 6.1 Rationale and Setup
State-of-the-art practice suggests using a reinforcement loop (Section 2.2) to improve code quality when using LLMs. Therefore, we constructed a feedback loop as a way to improve GPT-4o-generated code, starting by trying to eliminate the security issues identified in Section 5.2. For this experiment, we included 60 LeetCode files (solutions) generated by GPT-4o, as follows: 30 files with security issues (found by Clang), and 30 files without security issues. The list of security issues is listed in Table 7. We ask GPT-4o to regenerate these files based on different security categories. In each loop, prompts for files with security issues will be accompanied by prompts for files that have no security issues. For instance, we specifically ask GPT-4o to fix only malloc overflow issues in this loop with a total of 18 prompts related to files that have malloc overflow issues along with another 18 prompts related to files that have no malloc overflow issues. Note that the prompts (LeetCode problem descriptions) used in the loops are the same as the non-feedback setup, but with additional instructions that ask GPT-4o to make sure there are no certain security issues in the code. Since there are 3 categories of security issues, a total of 3 loops will be conducted, each loop to fix a specific security issue.
@@ 6.2 Results
Files with security issues. For files that did contain security issues, GPT-4o managed to lower the number of security issues found after one iteration on malloc overflow and null dereference from 40 to 24 and 3 to 2, respectively, as shown in Table 8. However, the security issues are not completely removed, even though we specifically asked GPT-4o to ensure the code will not have these security issues. Furthermore, for array index out of bounds issues, GPT-4o generated 3 more such issues. This is a critical finding since it shows the possibility that LLM may generate code that is the exact opposite of what users have specified in the prompt.